# Neural Architecture Search via Two Constant Shared Weights Initialisations

### v2

@@ 2.1 Empirical evaluation
2.1.1 NAS-Bench-201
The results for overall epsilon performance on NAS-Bench-201 are given in Table 1 along with other zero-cost NAS metrics. The Kendall $\tau$ score is not reported in Abdelfattah et al. (2021), but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation. We use the data provided by Abdelfattah et al. (2021) to evaluate their Kendal $\tau$. Note that our results differ from the original paper for some evaluation scores. In such cases, we indicate the original values between brackets. In particular, there is a discrepancy in computing the values in the last column, Top-64/top-$5\%$, while the rest of the results are consistent. Figure 6 in Appendix suggests that our calculations are correct.
Table 1: Zero-cost metrics performance for the NAS-Bench-201 search space with its three datasets: CIFAR-10, CIFAR-100 and ImageNet16-120. We give the original values from Abdelfattah et al. (2021) for reference between brackets. We highlight the best-performing metrics in bold.
@@ 2.1.1 NAS-Bench-201
epsilon 0.85 0.53 0.67 0.37 71.51 59
For NAS-Bench-201, we also report average performance when selecting one architecture from a pool of $N$ random architectures. The statistics are reported over $500$ runs. Table 2 compares epsilon to other trainless metrics. Note that te-nas starts with a super-network composed of all the edges and operators available within the space. In this case, $N$ is not applicable, and the performance can not be improved. In principle, other methods’ performance should improve with higher $N$ values.
Table 2: Comparison of the trainless metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. On the top, we list the best-performing methods that require training: REA (Real et al. (2019)), random search, REINFORCE (Williams (1992)), BOHB (Falkner et al. (2018)), DARTS- (Chu et al. (2020)). We report the average best-achieved test accuracy over $500$ runs, with $1{,}000$ architectures ($100$ for grad_sign) sampled from the search space at random. For tenas, the results are reported for $4$ random seeds. Random and optimal performances are given as a baseline.
@@
epsilon $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
Trainless (N=100)
@@
epsilon $0.003$ $\boldsymbol{90.44}\pm\boldsymbol{0.97}$ $\boldsymbol{93.39}\pm\boldsymbol{0.82}$ $\boldsymbol{70.85}\pm\boldsymbol{1.30}$ $\boldsymbol{71.00}\pm\boldsymbol{1.26}$ $\boldsymbol{44.03}\pm\boldsymbol{2.02}$ $\boldsymbol{44.20}\pm\boldsymbol{2.04}$
Trainless (all space)
tenas $0.5$ $93.9\pm 0.47$ $71.24\pm 0.56$ $42.38\pm 0.46$
Comparing the results for epsilon with other zero-cost NAS metrics, we can see that it outperforms them by a good margin. Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar figures for other methods can be found in Appendix, Figure 6). However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsilon development.
@@
Figure 1: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).
2.1.2 NAS-Bench-101
We use NAS-Bench-101 space to confirm that the success of epsilon metric in the previous section is not due to overfitting the NAS-Bench-201 search space and to see how it applies to a vaster search space. Table 3 together with Figure 2 confirm that it performs reasonably well on NAS-Bench-101, too.
Table 3: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. Values from Abdelfattah et al. (2021) are given for reference between brackets. We highlight the best-performing metrics in bold.
@@ 2.1.2 NAS-Bench-101
Figure 2: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure shows $423{,}624$ and $14{,}322$ networks for NAS-Bench-101 and NAS-Bench-NLP, respectively (excluding architectures with NaN scores).
2.1.3 NAS-Bench-NLP
Both NAS-Bench-201 and NAS-Bench-101 are created to facilitate NAS in image recognition. They operate convolutional networks of very similar constitutions. To truly probe the generalisability of the epsilon metric, we test it on NAS-Bench-NLP. Both input data format and architecture type differ from the first two search spaces.
Unfortunately, Abdelfattah et al. (2021) provides no data for NAS-Bench-NLP, disabling us from using their results for calculations. Therefore, in Table 4, we give only values provided in the paper together with our epsilon metric (data for ficher is absent). We want to note that, unlike accuracy, perplexity used for language-related ML problems should be minimised. Therefore, the signs of correlations with scoring metrics should be reversed, which is not the case for numbers given in Abdelfattah et al. (2021).
Table 4: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We highlight the best-performing metrics in bold.
@@ 2.1.3 NAS-Bench-NLP
epsilon -0.34 -0.12 -0.23 -0.08 24.87 11
@@ 2.2 Integration with other NAS methods
While it is possible to utilise zero-cost metrics independently, they are often implemented within other NAS. Here we provide examples of random search and ageing evolution algorithms when used in tandem with epsilon.
Similarly to Abdelfattah et al. (2021), we compare random search performance with and without warm-up. First, we create a warm-up pool of $3{,}000$ architectures. Then, during the first $64$ steps of random search, the algorithm picks networks from the pool based on the highest trainless score and accordingly updates the best test performance.
For ageing evolution, the same principle applies. En plus, we report the results of implementation where every next parent is decided based on the highest epsilon score (”move”). In other words, the trainless scoring metric replaces validation accuracy in move mode. Finally, the child is created by parent mutation (within an edit distance of $1$) and added to the pool. Finally, the oldest network is removed from the pool.
For both described algorithms, we run the procedure until the number of trained architectures reaches $300$ and perform $100$ random rounds. Figure 3 shows that epsilon metric leads to considerable improvements in terms of time and precision. The best performance is achieved in combination with a warm-up. Figure 5 assembles warm-up performances for several trainless metrics.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: epsilon integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
@@ 3 Ablation studies
While our metric is relatively straightforward to implement, it does require several hyperparameters to be set. In this section, we provide the ablations studies’ results and analyse them.
@@ 3.1 Weights
To compute the epsilon metric, we initialise the networks with two distinct constant shared weights. The question is: how to fix their values, and how does this choice affect the whole method?
To answer this question, we ran a series of tests with multiple pairs of weights. In pairs, it makes sense to set the first weight always greater than the second one. This consideration stems from the fact that epsilon is based on the mean absolute difference between two initialisations. Therefore, it will output the same score for [$w_{1}$, $w_{2}$] and [$w_{2}$, $w_{1}$] combinations and will be exactly zero in the case where the two weights are equal. Tables 5, 6, 7 summarises the results of our tests. For the NAS-Bench-201 search space, the outcomes are very similar between the datasets, and the minor differences do not affect the conclusions drawn from this data.
Table 5: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
@@
Table 6: Zero-cost metrics performance evaluated on NAS-Bench-201 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
@@
Table 7: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
@@
The results show that there is indeed some variation between different initialisations. Two effects impact performance. Firstly, the more significant the difference between the two weights, the higher the correlation between epsilon and trained performance. It makes sense since if the weights are too close to each other, the difference between the outputs becomes minimal, and the distinction between architectures in terms of epsilon is no longer possible.
Secondly, we can see that too low and too high weights result in fewer architectures with non-NaN scores. Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsilon metric. Naturally, the deeper the network is, the more signal attenuation is, and the higher the probability of NaN score. We remind that prior to epsilon computation, the outputs are normalised to the same range $[0,1]$, which means the value of the metric per se does not become smaller for deeper networks. This extinction effect is due to the limited sensitivity of float type. Theoretically, with infinite float sensitivity, there would be no extinction and hence no effects on the choice of the weights.
Table 8: Optimal weights for epsilon evaluation for three search spaces and their datasets.
@@
As a rule of thumb, for every new ML problem, we suggest running epsilon evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs. While we acknowledge that this procedure is subjective, importantly, it does not require any training.
@@ 3.2 Initialisation algorithm
The choice to initialise the weights with constant values provides a ground for discussion. On the one hand, constant shared weights guarantee the same initialisation regardless of the structure; the noise from the randomness of other initialisations is not obstructing the effect of the neural geometry. It would be possible to reduce the amount of noise by increasing the number of independent initialisations with distinct random seeds similar to Gracheva (2021), but this increases the algorithm run-time proportionally.
Table 9 gives the results of the ablation studies when we compute epsilon based on two initialisations.
Table 9: Initialisations ablation studies. The metric is computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
Initialisation Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
Uniform 15284 0.14 -0.17 0.09 -0.11 3.07 7
Normal 15284 0.44 -0.29 0.30 -0.20 20.07 4
Kaiming uniform 15284 0.32 0.11 0.21 0.07 27.40 26
Kaiming normal 15284 0.49 -0.29 0.33 -0.20 17.86 7
Orthogonal 15284 0.41 -0.20 0.28 -0.13 15.11 1
@@ 3.3 Batch size
To test the method’s sensitivity against the batch size, we ran the scoring routine with $8$ different batch sizes $[8,16,32,64,128,512,1024]$, each with $10$ random batches. The tests are done on the NAS-Bench-201 search space with its three datasets. For each batch, we report medians with $25$ and $75$ percentiles over $10$ runs. We evaluate epsilon on $5{,}000$ architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Figure 4 show that, as expected, the larger the batch size, the better and more stable the performance of epsilon metric. There is almost no improvement for batches over $256$. We set the batch size to this value in the experiments throughout the paper.

### v3

@@ 4.2 Evaluation metrics
5 Results
5.1 NAS-Bench-201
The results for overall epsinas performance on NAS-Bench-201 are given in Table 2 along with other zero-cost NAS metrics. The Kendall $\tau$ score is not reported in [1], but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation. We use the data provided by [1] to evaluate their Kendal $\tau$. Note that our results differ from the original paper in terms of evaluation scores. In such cases, we indicate the original values between brackets. In particular, there is a discrepancy in computing the values in the last column, Top-64/top-$5\%$, while the rest of the results are consistent. Figure 8 in the Appendix suggests that our calculations are correct.
Table 2: Zero-cost metrics performance for the NAS-Bench-201 search space with its three datasets: CIFAR-10, CIFAR-100 and ImageNet16-120. We give the original values from [1] for reference between brackets. We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
@@ 5.1 NAS-Bench-201
nparams 0.69 0.52
zico 0.79 0.60
$\xi$-gsnr 0.793 0.608
epsinas 0.85 0.53 0.67 0.37 71.51 59
For NAS-Bench-201, we also report average performance when selecting one architecture from a pool of $N$ random architectures. The statistics are reported over $500$ runs. Table 3 compares epsinas to other zero-cost metrics.
Table 3: Comparison of the zero-cost metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. On the top, we list the best-performing methods that require training: REA [38], random search, REINFORCE [49], BOHB [13], DARTS- [7]. We report the average best-achieved test accuracy over $500$ runs, with $1{,}000$ architectures ($100$ for grad_sign) sampled from the search space at random. Random and optimal performances are given as a baseline.
@@
epsinas $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
Zero-cost (N=100)
@@
epsinas $0.003$ $90.44\pm 0.97$ $93.39\pm 0.82$ $70.85\pm 1.30$ $71.00\pm 1.26$ $44.03\pm 2.02$ $44.20\pm 2.04$
Comparing the results for epsinas with other zero-cost NAS metrics, we can see that it shows surprisingly good performance given its conceptual simplicity. Importantly, it is consistently better than the naive estimator nparams.
Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar plots for other methods can be found in the Ablation section, Figure 8). However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsinas development.
@@
Figure 1: Zero-cost NAS epsinas metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).
5.2 NAS-Bench-101
We use the NAS-Bench-101 search space to verify that the effectiveness of the epsinas metric, as demonstrated in the previous section, is not merely a result of overfitting to the NAS-Bench-201 space. It also allows us to assess its performance in a broader search space. As shown in Table 4 and Figure 2, epsinas performs well on NAS-Bench-101, once again surpassing the nparams proxy.
Table 4: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. Values from [1] are given for reference between brackets. We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
@@ 5.2 NAS-Bench-101
Figure 2: Zero-cost NAS epsinas metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure shows $423{,}624$ and $14{,}322$ networks for NAS-Bench-101 and NAS-Bench-NLP, respectively (excluding architectures with NaN scores).
5.3 NAS-Bench-NLP
Both NAS-Bench-201 and NAS-Bench-101 are created to facilitate NAS in image recognition. They operate convolutional networks of very similar constitutions. To truly probe the generalisability of the epsinas metric, we test it on NAS-Bench-NLP. Both input data format and architecture type differ from the first two search spaces.
Unfortunately, [1] provides no data for NAS-Bench-NLP, disabling us from using their results for calculations. Therefore, in Table 5, we give only values provided in the paper together with our epsinas metric (data for ficher is absent). It’s important to note that, unlike accuracy, perplexity – commonly used in language-related machine learning problems – should be minimized. Thus, the signs of correlations with scoring metrics must be reversed, which differs from the results reported in [1].
Table 5: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We highlight the best-performing metrics in bold (p-values between epsinas and other metrics are below 0.001).
@@ 5.3 NAS-Bench-NLP
epsinas -0.34 -0.12 -0.24 -0.08 26.23 13
5.4 Computational cost
One of the primary purposes of zero-shot proxies is to reduce the computational cost of the NAS process. Since epsinas only requires two forward passes through the data and does not involve gradient computation, it is efficient regarding both time and memory usage. Figure 3 presents the results of tests conducted using NAS-Bench-101. As anticipated, memory usage and processing time scale linearly with the number of parameters and FLOPs. It also demonstrates that epsinas can be computed using a CPU within 30 seconds for the largest NAS-Bench-101 networks consisting of several million parameters.
Figure 3: Computational efficiency of the zero-cost NAS epsinas metric for the NAS-Bench-101 search space. The results are based on the evaluation of $10{,}000$ networks from the NAS-Bench-101 benchmark, divided into $20$ bins according to the number of parameters. The shaded area represents one standard deviation within each group. FLOPs are reported for a batch size of $1$, while memory usage and processing time are evaluated with a batch size of $256$.
Comparing epsinas to other zero-cost metrics is not straightforward, as different methods use different search spaces, computational setups, and batch sizes. For instance, zico reports evaluation times using MobileNetV2 [41] with a budget of $1,000$ MFLOPs and a batch size of $128$. The authors noted that evaluating $10^{5}$ architectures required $10$ GPU hours.
MobileNetV2 networks range from $1.7$M to $6.9$M parameters, larger than NAS-Bench-201 but smaller than NAS-Bench-101. To provide a closer comparison, we evaluated epsinas on NAS-Bench-101 with a batch size of $128$, limiting architectures to $6.9$M parameters. Under these conditions, evaluating $10^{5}$ architectures took approximately $7$ GPU hours, with the epsinas computation requiring only $2$ GPU hours; the remaining time was spent on model loading.
While this is not a direct comparison, epsinas does not involve gradient computation, so it is reasonable to expect it to be faster per architecture evaluation. The results support this expectation.
5.5 Correlation with nparams
As noted in previous studies [35, 25, 48], one of the best naive evaluators of a network’s performance is the number of parameters, nparams. Generally, the more parameters a network has, the better we expect its predictive performance to be. However, bigger does not always mean better: large networks are susceptible to overfitting and can be prohibitively expensive to train [8, 22]. Moreover, research on network pruning has shown that, with the right architecture, smaller subnetworks can achieve comparable or even superior performance to their larger counterparts [15]. Hence, when seeking an optimal evaluation proxy, it is essential to ensure that it can effectively identify smaller networks. Figure 4 illustrates the relationship between epsinas and the number of parameters, showing no strong preference for larger network sizes.
Figure 4: Visualisation of the relationship between epsinas score and the number of parameters in a network, nparams for CIFAR-10 data of NAS-Bench-201 benchmark.
@@ 5.6 Integration with other NAS methods
While zero-cost metrics are robust on their own, they benefit from being integrated with other NAS algorithms. [32] showed that combining a zero-cost metric based on Jacobian covariance, similar to [34], with REA [38] speeds up evolutionary search and helps find top-performing architectures. This work demonstrates how the epsinas metric can improve random search and ageing evolution algorithms.
This work follows the implementation of [1].
Random search Each new architecture is selected randomly and evaluated based on test accuracy. The final test performance is reported without using any auxiliary metrics.
Random search with warm-up A warm-up pool of $3{,}000$ randomly sampled architectures is first ranked using epsinas. Next, models are ”trained” in order of decreasing epsinas value by retrieving their test accuracies. Training stops once the budget of $300$ architectures is reached.
Ageing evolution We implement the basic ageing evolution algorithm from [38]. The process starts with a randomly sampled pool of $64$ architectures. Parents are selected based on test accuracy, and a child is generated through mutation (with an edit distance of $1$). The child is added to the pool, while the oldest architecture is removed. This process repeats until $300$ architectures have been trained.
Ageing evolution with warm-up A warm-up phase begins with $3{,}000$ sampled architectures, which are ranked using the epsinas metric. The $64$ architectures with the highest epsinas values form the initial pool. The standard ageing evolution process continues, where parents are selected based on test accuracy.
Ageing evolution with move Instead of using test accuracy for parent selection, architectures are assessed based on their epsinas scores.
We conduct $100$ random runs for each algorithm and report the mean and standard deviation.
Figure 5 demonstrates that the epsinas metric significantly improves both efficiency and accuracy, with the best results achieved when combined with a warm-up. Figure 7 compares warm-up performances across various zero-cost metrics.
Figure 5: epsinas integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
@@ 6 Ablation studies
While our metric is straightforward to implement, it depends on several hyperparameters. In this section, we present and analyse the results of our ablation studies.
@@ 6.1 Weights
To compute the epsinas metric, we initialise the networks with two distinct constant shared weights. The question is: how do we fix their values, and how does this choice affect the whole method?
To answer this question, we ran a series of tests with multiple pairs of weights. It makes sense to set the first weight always greater than the second one in pairs. This consideration stems from the fact that epsinas is based on the mean absolute difference between two initialisations. Therefore, it will output the same score for [$w_{1}$, $w_{2}$] and [$w_{2}$, $w_{1}$] combinations and will be exactly zero in the case where the two weights are equal. Tables 6, 7, 8 summarise the results of our tests. For the NAS-Bench-201 search space, the outcomes are very similar between the datasets, and the minor differences do not affect the conclusions drawn from this data.
Table 6: Zero-cost metrics performance was evaluated using the NAS-Bench-101 search space and the CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values is given in the second column.
@@
Table 7: Zero-cost metrics performance was evaluated in the NAS-Bench-201 search space and the CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values is given in the second column.
@@
Table 8: Zero-cost metrics performance was evaluated using the NAS-Bench-NLP search space and the PTB dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values is given in the second column.
@@
The results indicate that variations in initialisation do affect performance. Two key factors influence this impact. First, a more significant difference between the two weights leads to a higher correlation between epsinas and trained performance. It is intuitive – when the weights are too similar, the output difference becomes minimal, making it difficult to distinguish between architectures based on epsinas.
Secondly, we can see that too low and too high weights result in fewer architectures with non-NaN scores. Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsinas metric. Naturally, the deeper the network is, the more signal attenuation is, and the higher the probability of NaN score. We remember that before epsinas computation, the outputs are normalised to the same range $[0,1]$, which means the metric value per se does not become smaller for deeper networks. This extinction effect is due to the limited sensitivity of float type. Theoretically, with infinite float sensitivity, there would be no extinction and hence no effects on the choice of the weights.
Table 9: Optimal weights for epsinas evaluation for three search spaces and their datasets.
@@
As a rule of thumb, for every new ML problem, we suggest running epsinas evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs. While we acknowledge that this procedure is subjective, importantly, it does not require any training.
@@ 6.2 Initialisation algorithm
The choice to initialise the weights with constant values provides a ground for discussion. On one hand, using constant shared weights ensures a consistent initialisation across different structures, eliminating the noise introduced by random initialisations. It allows us to isolate and analyse the impact of neural geometry more effectively. Averaging results over multiple runs with different random seeds, as in [18], can reduce the randomness noise but increases computational time proportionally.
Table 10 presents the results of our ablation studies, where epsinas is computed using different initialisation schemes. The correlation coefficients are based on the epsinas scores from two initialisations but using standard initialisation methods instead of constant shared weights. For every initialisation scheme, we only modify the convolutional layers. The rest of the architecture is initialised as per the original benchmark space: constant initialisation with ones for the batch normalisation, normal uniform for the linear layer, and biases set to zero.
Table 10: Initialisations ablation studies. The metric is computed over the first $5{,}000$ architectures in the NAS-Bench-201 benchmark.
Initialisation Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
@@
Uniform -0.02 -0.01 -0.01 -0.00 7.13 13
Normal 0.08 -0.20 0.05 -0.13 15.83 4
Kaiming uniform 0.07 -0.17 0.05 0.12 13.60 2
Kaiming normal 0.07 -0.13 0.05 -0.09 12.49 1
Orthogonal 0.07 -0.16 0.05 -0.10 11.71 6
@@ 6.3 Batch size
To test the method’s sensitivity against the batch size, we ran the scoring routine with $8$ different batch sizes $[8,16,32,64,128,512,1024]$, each with $10$ random batches. The tests are done on the NAS-Bench-201 search space with its three datasets. We report medians with $25$ and $75$ percentiles over $10$ runs for each batch. We evaluate epsinas on $5{,}000$ architectures.
@@
Figure 6 shows that, as expected, the larger the batch size, the better and more stable the performance of epsinas metric. There is almost no improvement for batches over $256$. We set the batch size to this value in the experiments throughout the paper.